# Adapted optimal transport between Gaussian processes in discrete time

### v1

@@ Abstract.
We derive explicitly the adapted $2$-Wasserstein distance between arbitrary non-degenerate Gaussian distributions on $\mathbb{R}^{N}$ and characterize the optimal bicausal coupling(s). This leads to an adapted version of the Bures-Wasserstein distance on the space of positive definite matrices.
@@
Adapted optimal transport, Wasserstein distance, Bures-Wasserstein, Knothe-Rosenblatt, Gaussian process
@@ 1. Introduction
Adapted (by which we mean bicausal) optimal transport (AOT) has emerged to be a useful tool for quantifying distributional uncertainty and the sensitivity of stochastic optimization problems in contexts where the flow of information in time plays a crucial role. The key idea, which has several origins in different fields, is to impose a (bi)causal constraint on the feasible couplings. We recall the necessary concepts in Section 2 and refer the reader to [3, 5, 4, 6, 9, 14, 18, 24], and the references therein, for detailed expositions of the theory and the related literature.
Several AOT problems have been explicitly solved. For example, [6] provides a sufficient condition for the Knothe-Rosenblatt coupling (recalled in Section 3) to be optimal in discrete time. Also see [7, 12, 18, 25] for continuous time results concerning Markovian stochastic differential equations. The main purpose of this paper is to address the adapted $2$-Wasserstein transport between arbitrary non-degenerate Gaussian measures on $\mathbb{R}^{N}$ which are possibly non-Markovian. To the best of our knowledge, this basic case is still open despite the rapid growth of the subject in recent years. The closest papers we could find are [16], which studies a distributionally robust filtering problem with Gaussian noise, and [27], which solves a related OT problem between stationary Gaussian processes.
@@
where $d_{\mathrm{ABW}}$ is the adapted Bures-Wasserstein distance on the space $\mathscr{S}_{++}(N)$ of positive definite matrices defined by
@@
Here diag()normal-$\operatorname{diag}(\cdot)$ gives the diagonal of a matrix, $\|\cdot\|_{1}$ is the $\ell_{1}$-norm and $\|\cdot\|=\|\cdot\|_{2}$ is the Euclidean norm.
After reviewing the Knothe-Rosenblatt coupling in Section 3, we prove Theorem 1.1 in Section 4 using a dynamic programming principle from [6] and characterize the optimal bicausal couplings in Corollary 4.4. Since the Gaussian distribution is fundamental in various applications, our explicit solution may improve the understanding of the general theory and stimulate new applications of AOT. We also note that Knothe-Rosenblatt transport maps (also called monotone triangular transport maps) have found recent applications in statistics and machine learning [2, 8]. Some geometric properties of the adapted Bures-Wasserstein distance are discussed in Section 5. Section 6 provides illustrative examples. Finally, we discuss some future directions in Section 7.
@@ 2. Adapted optimal transport
We work with Borel probability measures on $\mathbb{R}^{N}$, $N\in\mathbb{N}$, regarded as laws of real-valued stochastic processes indexed by $t\in[N]:=\{1,\ldots,N\}$. We use $\mathcal{P}(E)$ to denote the collection of Borel probability measures on a space $E$. By a coupling of $(\mu,\nu)\in(\mathcal{P}(\mathbb{R}^{N}))^{2}$ we mean an element $\pi\in\mathcal{P}(\mathbb{R}^{N}\times\mathbb{R}^{N})$ whose first and second marginals are $\mu$ and $\nu$ respectively. The set of couplings of $(\mu,\nu)$ is denoted by $\Pi(\mu,\nu)$. By an abuse of terminology we also use the word coupling to mean a pair $(X=(X_{t})_{t=1}^{T},Y=(Y_{t})_{t=1}^{N})$ of real-valued processes, defined on some probability space, with joint law $\pi\in\Pi(\mu,\nu)$. We write $\mathbb{P}_{\pi}$ for a probability under which $(X,Y)\sim\pi$ (analogous for $\mathbb{E}_{\pi}$).
@@
For integers $s\leq t$ we let $s:t=(s,s+1,\ldots,t)$. We also write $\mathbf{t}=1:t$ and $\mathbf{t}^{\prime}=(t+1):N$. For a quantity $Z$ indexed by time we use the shorthand $Z_{s:t}=(Z_{s},\ldots,Z_{t})$ (by convention $Z_{1:0}$ is empty). Suppose $\mu\in\mathcal{P}(\mathbb{R}^{N})$ and $X\sim\mu$. We let $\mu_{t}(\cdot)$ be the marginal distribution of $X_{t}$. For $x_{1:t}\in\mathbb{R}^{t}$, we let $\mu_{t+1}(\cdot|x_{1:t})\in\mathcal{P}(\mathbb{R})$ be the conditional distribution of $X_{t+1}$ given $X_{1:t}=x_{1:t}$, and let $\mu_{\mathbf{t}^{\prime}}(\cdot|x_{\mathbf{t}})$ be the conditional distribution of $X_{\mathbf{t}^{\prime}}$ given $X_{\mathbf{t}}=x_{\mathbf{t}}$. Analogously, we define $\nu_{t+1}(\cdot|y_{\mathbf{t}})$ and $\nu_{\mathbf{t}^{\prime}}(\cdot|y_{\mathbf{t}})$ for $Y$, as well as $\pi_{t+1}(\cdot|x_{\mathbf{t}},y_{\mathbf{t}})$ and $\pi_{{\bf t}^{\prime}}(\cdot|x_{\mathbf{t}},y_{\mathbf{t}})$ for $(X,Y)$. These (regular) conditional distributions are well-defined by the disintegration theorem and are a.s. unique with respect to suitable reference measures.
@@
Let $\mathcal{P}_{2}(\mathbb{R}^{N})=\{\mu\in\mathcal{P}(\mathbb{R}^{N}):\int_{\mathbb{R}^{N}}\|x\|^{2}\mu(\mathrm{d}x)<\infty\}$ be the collection of probability measures on $\mathbb{R}^{N}$ with finite second moment. For $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{N})$, the usual $2$-Wasserstein distance $\mathcal{W}_{2}(\mu,\nu)$ is defined by
(2.1) $\mathcal{W}_{2}^{2}(\mu,\nu)=\inf_{\pi\in\Pi(\mu,\nu)}\int_{\mathbb{R}^{N}\times\mathbb{R}^{N}}\|x-y\|^{2}\mathrm{d}\pi(x,y).$
@@
We simply call $\mathcal{W}_{2}$ the Wasserstein distance since $2$ is the only order considered in this paper (similar for other quantities such as the Knothe-Rosenblatt distance (3.2)). Let $\mathscr{S}_{++}(N)$ (resp. $\mathscr{S}_{+}(N)$) be the space of $N$-by-$N$ positive definite (resp. semipositive definite) matrices. When $\mu=\mathcal{N}(a,A)$ and $\nu=\mathcal{N}(b,B)$ are Gaussian distributions on $\mathbb{R}^{N}$, [^1] We write $\mathcal{N}_{N}(a,A)$ when there is a need to emphasize the dimension. it is well known that
(2.2) $\mathcal{W}_{2}^{2}(\mu,\nu)=\|a-b\|^{2}+d_{\mathrm{BW}}^{2}(A,B),$
where $d_{\mathrm{BW}}$ is the Bures-Wasserstein distance on $\mathscr{S}_{+}(N)$ defined by
@@
Here $A^{\frac{1}{2}}\in\mathscr{S}_{+}(N)$ is the unique matrix square root of $A\in\mathscr{S}_{+}(N)$. See [11, 20, 26] for in-depth studies of $d_{\mathrm{BW}}$ from the viewpoints of matrix analysis and Riemannian geometry. When $A,B\in\mathscr{S}_{++}(N)$ the value $\mathcal{W}_{2}^{2}(\mu,\nu)$ is achieved uniquely by the deterministic coupling $Y=T_{\mathrm{W}}^{\mu,\nu}X$, where $X\sim\mu$ and
(2.4) $T_{\mathrm{W}}^{\mu,\nu}x=b+A^{-\frac{1}{2}}\big{(}A^{\frac{1}{2}}BA^{\frac{1}{2}}\big{)}^{\frac{1}{2}}A^{-\frac{1}{2}}(x-a).$
Here and below we use the column convention for $X$, $Y$, etc. When $N=1$ this reduces to the comonotonic coupling with transport map $T_{\mathrm{W}}^{\mu,\nu}=F_{\nu}^{-1}\circ F_{\mu}$, where $F$ denotes the distribution function. For later use we also recall the counter-monotonic coupling which has transport map $F_{\nu}^{-1}\circ R\circ F_{\mu}$, where $R(u)=1-u$. Note that the Jacobian matrix in (2.4) is an element of $\mathscr{S}_{++}(N)$ since $T_{\mathrm{W}}^{\mu,\nu}$ is a convex gradient. Unless it is diagonal, the transport map “looks into the future of $X$” since generally each $Y_{t}$ is a function of both $X_{\mathbf{t}}$ and $X_{\mathbf{t}^{\prime}}$. The causal condition excludes such couplings.
@@
We cite a useful equivalent condition, taken from [6, Proposition 5.1], for the bicausal property. Given $\pi\in\Pi(\mu,\nu)$, we decompose it as a product of regular conditional distributions:
(2.5) $\begin{split}\pi(\mathrm{d}x_{1:N},\mathrm{d}y_{1:N})&=\pi_{1}(\mathrm{d}x_{1},\mathrm{d}y_{1})\pi_{2}(\mathrm{d}x_{2},\mathrm{d}y_{2}|x_{1},y_{1})\\ &\quad\quad\cdots\pi_{N}(\mathrm{d}x_{N},\mathrm{d}y_{N}|x_{1:(N-1)},y_{1:(N-1)}).\end{split}$
Then $\pi\in\Pi_{bc}(\mu,\nu)$ if and only if $\pi_{1}\in\Pi(\mu_{1},\nu_{1})$ and, for all $t$ and ($\pi$-almost) all $x_{1:t},y_{1:t}\in\mathbb{R}^{t}$, we have
$\pi_{t+1}(\cdot|x_{1:t},y_{1:t})\in\Pi(\mu_{t+1}(\cdot|x_{1:t}),\nu_{t+1}(\cdot|y_{1:t})).$
If in (2.1) we optimize over bicausal couplings we obtain the adapted Wasserstein distance.
@@
(2.6) $\mathcal{AW}_{2}^{2}(\mu,\nu)=\inf_{\pi\in\Pi_{bc}(\mu,\nu)}\int_{\mathbb{R}^{N}\times\mathbb{R}^{N}}\|x-y\|^{2}\mathrm{d}\pi(x,y).$
One can show that $\mathcal{AW}_{2}$ is a metric on $\mathcal{P}_{2}(\mathbb{R}^{N})$; its induced topology is the fundamental weak adapted topology which has several equivalent characterizations [3, 24]. In this paper we focus on the case where $\mu$ and $\nu$ are Gaussian, i.e., $X$ and $Y$ are Gaussian processes.
The next lemma is standard and the proof is omitted.
Lemma 2.3 (Centering).
Let $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{N})$. Let $a=\int_{\mathbb{R}^{N}}x\mu(\mathrm{d}x)$ and $b=\int_{\mathbb{R}^{N}}y\nu(\mathrm{d}y)$ be the means $\mu$ and $\nu$ respectively. Then $\mathcal{AW}_{2}^{2}(\mu,\nu)=\|a-b\|^{2}+\mathcal{AW}_{2}^{2}(\tilde{\mu},\tilde{\nu})$, where $\tilde{\mu}=(\cdot-a)_{\#}\mu$ and $\tilde{\nu}=(\cdot-b)_{\#}\nu$ are the centered versions of $\mu$ and $\nu$. The same result holds for $\mathcal{W}_{2}$.
@@ 3. The Knothe-Rosenblatt coupling
For $\mu,\nu\in\mathcal{P}(\mathbb{R}^{N})$, the Knothe-Rosenblatt coupling, also called the multivariate quantile transform, defines a bicausal coupling which is known to be $\mathcal{AW}_{2}$-optimal under a monotonicity condition on $\mu$ and $\nu$ (see [6, Proposition 5.3]). We consider directly the Gaussian case where $\mu=\mathcal{N}(a,A)$, $\nu=\mathcal{N}(b,B)$ and $A,B\in\mathscr{S}_{++}(N)$. It is useful to parameterize $A$ and $B$ in terms of their Cholesky decompositions. Let $\mathscr{L}_{++}(N)$ (resp. $\mathscr{L}_{+}(N)$) be the space of $N$-by-$N$ lower triangular matrices with positive (resp. non-negative) diagonal entries. Then there exist unique $L,M\in\mathscr{L}_{++}(N)$ such that $A=LL^{\top}$ and $B=MM^{\top}$. It is not difficult to show (see, e.g., [19, Proposition 2]) that the Cholesky map $A\mapsto L$ is a diffeomorphism from $\mathscr{S}_{++}(N)$ onto $\mathscr{L}_{++}(N)$. When $A$ is only positive semidefinite there exists $L\in\mathscr{L}_{+}(N)$ such that $A=LL^{\top}$ but $L$ is not necessarily unique (an example is given in (5.1)). On some probability space let $\epsilon\sim\mathcal{N}(0,I)$ be a standard Gaussian random vector in $\mathbb{R}^{N}$. The Knothe-Rosenblatt coupling $\pi_{\mathrm{KR}}^{\mu,\nu}$ between $\mu$ and $\nu$ is given by the law of $(X,Y)=(a+L\epsilon,b+M\epsilon)$. In this context, we may also call $\pi_{\mathrm{KR}}^{\mu,\nu}$ the synchronous coupling since the same noise sequence $\epsilon$ is used to drive both $X$ and $Y$ regarded as solutions to suitable stochastic difference equations. It is clear that $\pi_{\mathrm{KR}}^{\mu,\nu}$ is a deterministic coupling under which
(3.1) $Y=T_{\mathrm{KR}}^{\mu,\nu}X,\quad T_{\mathrm{KR}}^{\mu,\nu}x=b+ML^{-1}(x-a).$
Note that by construction we have $ML^{-1}\in\mathscr{L}_{++}(N)$.
Following [10], we define
(3.2) $\mathcal{KR}_{2}(\mu,\nu)=\left(\int_{\mathbb{R}^{N}\times\mathbb{R}^{N}}\|x-y\|^{2}\mathrm{d}\pi_{\mathrm{KR}}^{\mu,\nu}(x,y)\right)^{1/2}$
and call it the Knothe-Rosenblatt distance between $\mu$ and $\nu$.
@@
For $\mu=\mathcal{N}(a,A)$ and $\nu=\mathcal{N}(b,B)$ we have
(3.3) $\mathcal{KR}_{2}^{2}(\mu,\nu)=\|a-b\|^{2}+d_{\mathrm{KR}}^{2}(A,B),$
where $d_{\mathrm{KR}}$ is the Knothe-Rosenblatt distance on $\mathscr{S}_{++}(N)$ defined by
(3.4) $d_{\mathrm{KR}}^{2}(A,B)=\operatorname{tr}(A)+\operatorname{tr}(B)-2\operatorname{tr}(L^{\top}M)=\|L-M\|_{\mathrm{F}}^{2},$
and $\|L\|_{\mathrm{F}}=\sqrt{\operatorname{tr}(L^{\top}L)}$ is the Frobenius norm.

### v4

@@ Abstract.
We derive explicitly the adapted $2$-Wasserstein distance between non-degenerate Gaussian distributions on $\mathbb{R}^{N}$ and characterize the optimal bicausal coupling(s). This leads to an adapted version of the Bures-Wasserstein distance on the space of positive definite matrices.
@@
Adapted optimal transport, Wasserstein distance, Bures-Wasserstein, Knothe-Rosenblatt, adapted Brenier, Gaussian process
@@ 1. Introduction
Adapted optimal transport (AOT) has emerged to be an appropriate framework for quantifying distributional uncertainty and the sensitivity of stochastic optimization problems in contexts where the flow of information in time plays a crucial role. The key idea, which has several origins in different fields, is to impose a (bi)causal constraint on the feasible couplings. We recall the necessary concepts in Section 2 and refer the reader to [5, 6, 7, 8, 11, 13, 18, 20, 26], and the references therein, for detailed expositions of the theory and the related literature.
Several AOT problems have been explicitly solved. For example, [8, 9, 28] provide sufficient conditions for the Knothe-Rosenblatt coupling, recalled in Section 3, to be optimal in discrete time. Also see [9, 15, 20, 27] for continuous-time results concerning stochastic differential equations. The main purpose of this paper is to address the adapted $2$-Wasserstein transport between arbitrary non-degenerate Gaussian measures on $\mathbb{R}^{N}$ which are possibly non-Markovian. To the best of our knowledge, this basic case is still open despite the rapid growth of the subject in recent years. The closest papers we could find are [19], which studies a distributionally robust filtering problem with Gaussian noise, and [32], which solves a related OT problem between stationary Gaussian processes. Now we state the main theorem of the paper.
@@
where $d_{\mathrm{ABW}}$ is the adapted Bures-Wasserstein distance on the space $\mathscr{S}_{++}(N)$ of (symmetric) positive definite matrices defined by
@@
Here $\operatorname{diag}(\cdot)$ gives the diagonal of a square matrix, $\|\cdot\|_{1}$ is the $\ell_{1}$-norm and $\|\cdot\|=\|\cdot\|_{2}$ is the Euclidean norm. (See Remark 4.5(i) for an extension to a weighted square cost.)
In Section 4 we prove Theorem 1.1 via a dynamic programming principle from [8] and characterize the set of optimal bicausal couplings (Corollary 4.4). The main idea is to observe that the Gaussian assumption and the sum separability of the quadratic cost lead to a tractable linear-quadratic structure of the value function (see (4.9)). Recently, the authors of [1] built on our techniques to extend Theorem 1.1 to multivariate Gaussian processes together with entropic regularization. We discuss some geometric properties of the adapted Bures-Wasserstein distance in Section 5, and provide illustrative examples in Section 6. Since the Gaussian distribution is fundamental in various applications, our explicit solution and its extensions in [1] may improve the understanding of the general theory and stimulate new applications of AOT. Some directions for future research are discussed in Section 7.
@@ 2. Adapted optimal transport
We work with Borel probability measures on $\mathbb{R}^{N}$, $N\geq 1$, regarded as laws of real-valued stochastic processes indexed by $t\in[N]:=\{1,\ldots,N\}$ and equipped with their natural filtrations (implicit in (2.4) below). We use $\mathcal{P}(E)$ to denote the collection of Borel probability measures on a space $E$. By a coupling of $(\mu,\nu)\in(\mathcal{P}(\mathbb{R}^{N}))^{2}$ we mean an element $\pi\in\mathcal{P}(\mathbb{R}^{N}\times\mathbb{R}^{N})$ whose first and second marginals are $\mu$ and $\nu$ respectively. The set of couplings of $(\mu,\nu)$ is denoted by $\Pi(\mu,\nu)$. By an abuse of terminology we also use the word coupling to mean a pair $(X=(X_{t})_{t=1}^{N},Y=(Y_{t})_{t=1}^{N})$ of real-valued processes, defined on some probability space, with joint law $\pi\in\Pi(\mu,\nu)$. We write $\mathbb{P}_{\pi}$ for a probability under which $(X,Y)\sim\pi$ (analogous for the expectation $\mathbb{E}_{\pi}$).
@@
For integers $s\leq t$ we let $s:t=(s,s+1,\ldots,t)$. We also write $\mathbf{t}=1:t$ and $\mathbf{t}^{\prime}=(t+1):N$. For a quantity $Z$ indexed by $t\in[N]$ we use the shorthand $Z_{s:t}=(Z_{s},\ldots,Z_{t})$ (by convention $Z_{1:0}$ and $Z_{(N+1):N}$ are empty). Suppose $\mu\in\mathcal{P}(\mathbb{R}^{N})$ and $X\sim\mu$. For $t\in[N]$ we let $\mu_{t}(\cdot)\in\mathcal{P}(\mathbb{R})$ be the marginal distribution of $X_{t}$. For $x_{1:t}\in\mathbb{R}^{t}$, we let $\mu_{t+1}(\cdot|x_{1:t})\in\mathcal{P}(\mathbb{R})$ be the conditional distribution of $X_{t+1}$ given $X_{1:t}=x_{1:t}$, and let $\mu_{\mathbf{t}^{\prime}}(\cdot|x_{\mathbf{t}})\in\mathcal{P}(\mathbb{R}^{N-t})$ be the conditional distribution of $X_{\mathbf{t}^{\prime}}$ given $X_{\mathbf{t}}=x_{\mathbf{t}}$. Analogously, we define $\nu_{t+1}(\cdot|y_{\mathbf{t}})$ and $\nu_{\mathbf{t}^{\prime}}(\cdot|y_{\mathbf{t}})$ for $Y$, as well as $\pi_{t+1}(\cdot|x_{\mathbf{t}},y_{\mathbf{t}})$ and $\pi_{{\bf t}^{\prime}}(\cdot|x_{\mathbf{t}},y_{\mathbf{t}})$ for $(X,Y)$. These (regular) conditional distributions are well-defined by the disintegration theorem and are a.s. unique with respect to suitable reference measures.
@@
Let $\mathcal{P}_{2}(\mathbb{R}^{N})=\{\mu\in\mathcal{P}(\mathbb{R}^{N}):\int_{\mathbb{R}^{N}}\|x\|^{2}\mu(\mathrm{d}x)<\infty\}$ be the collection of Borel probability measures on $\mathbb{R}^{N}$ with finite second moment. For $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{N})$, the usual $2$-Wasserstein distance $\mathcal{W}_{2}(\mu,\nu)$ is defined by
(2.1) $\mathcal{W}_{2}^{2}(\mu,\nu)=\inf_{\pi\in\Pi(\mu,\nu)}\int_{\mathbb{R}^{N}\times\mathbb{R}^{N}}\|x-y\|^{2}\pi(\mathrm{d}x\mathrm{d}y).$
@@
We simply call $\mathcal{W}_{2}$ the Wasserstein distance since $2$ is the only order considered in this paper (similar for other quantities such as the Knothe-Rosenblatt distance (3.3)). When $\mu$ is absolutely continuous, Brenier’s theorem asserts that there exists a unique optimal coupling $\pi_{\mathrm{W}}^{\mu,\nu}$ given in terms of the gradient of a convex function [30]. Let $\mathscr{S}_{++}(N)$ (resp. $\mathscr{S}_{+}(N)$) be the space of $N$-by-$N$ (symmetric) positive definite (resp. semipositive definite) matrices. When $\mu=\mathcal{N}(a,A)$ and $\nu=\mathcal{N}(b,B)$ are Gaussian distributions on $\mathbb{R}^{N}$, [^1] We write $\mathcal{N}_{N}(a,A)$ when there is a need to emphasize the dimension. it is well known that $\mathcal{W}_{2}^{2}(\mu,\nu)=\|a-b\|^{2}+d_{\mathrm{BW}}^{2}(A,B)$, where $d_{\mathrm{BW}}$ is the Bures-Wasserstein distance on $\mathscr{S}_{+}(N)$ defined by
@@
Here $A^{\frac{1}{2}}\in\mathscr{S}_{+}(N)$ is the unique matrix square root of $A\in\mathscr{S}_{+}(N)$. See [14, 22, 29] for in-depth studies of $d_{\mathrm{BW}}$ from the viewpoints of matrix analysis and Riemannian geometry. When $A,B\in\mathscr{S}_{++}(N)$, $\pi_{\mathrm{W}}^{\mu,\nu}$ is the deterministic coupling under which $Y=T_{\mathrm{W}}^{\mu,\nu}X$, where $X\sim\mu$ and
(2.3) $T_{\mathrm{W}}^{\mu,\nu}x=b+A^{-\frac{1}{2}}\big{(}A^{\frac{1}{2}}BA^{\frac{1}{2}}\big{)}^{\frac{1}{2}}A^{-\frac{1}{2}}(x-a),\quad x\in\mathbb{R}^{N},$
is the Brenier map from $\mu$ to $\nu$. Here and below we use the column convention for $X$, $Y$, $a$, $b$, etc. When $N=1$ this reduces to the comonotonic coupling induced by the non-decreasing transport map $F_{\nu}^{-1}\circ F_{\mu}$, where $F$ denotes the distribution function. For later use, we also recall the counter-monotonic coupling which has the non-increasing transport map $F_{\nu}^{-1}\circ R\circ F_{\mu}$, where $R(u)=1-u$. Note that the Jacobian matrix in (2.3) is an element of $\mathscr{S}_{++}(N)$ since $T_{\mathrm{W}}^{\mu,\nu}$ is a convex gradient. Unless it is diagonal, the transport map $T_{\mathrm{W}}^{\mu,\nu}$ “looks into the future of $X$” since generally each $Y_{t}$ is a function of both $X_{\mathbf{t}}$ and $X_{\mathbf{t}^{\prime}}$. The causal condition excludes such couplings.
@@
We cite an equivalent condition, taken from [8, Proposition 5.1], for the bicausal property. Given $\pi\in\Pi(\mu,\nu)$, decompose it as a product of regular conditional distributions:
(2.5) $\begin{split}\pi(\mathrm{d}x\mathrm{d}y)=\pi_{1}(\mathrm{d}x_{1}\mathrm{d}y_{1})\pi_{2}(\mathrm{d}x_{2}\mathrm{d}y_{2}|x_{1},y_{1})\cdots\pi_{N}(\mathrm{d}x_{N}\mathrm{d}y_{N}|x_{1:(N-1)},y_{1:(N-1)}).\end{split}$
Then $\pi\in\Pi_{bc}(\mu,\nu)$ if and only if $\pi_{1}\in\Pi(\mu_{1},\nu_{1})$ and, for all $11\leq t\leq N-$ 1 and ($\pi$-almost) all $x_{1:t},y_{1:t}\in\mathbb{R}^{t}$, we have $\pi_{t+1}(\cdot|x_{1:t},y_{1:t})\in\Pi(\mu_{t+1}(\cdot|x_{1:t}),\nu_{t+1}(\cdot|y_{1:t}))$. In [28] this is called a Markov-construction. If in (2.1) we optimize over bicausal couplings, we obtain the adapted Wasserstein distance.
@@
(2.6) $\mathcal{AW}_{2}^{2}(\mu,\nu)=\inf_{\pi\in\Pi_{bc}(\mu,\nu)}\int_{\mathbb{R}^{N}\times\mathbb{R}^{N}}\|x-y\|^{2}\pi(\mathrm{d}x\mathrm{d}y).$
Clearly $\mathcal{AW}_{2}\geq\mathcal{W}_{2}$, and equality holds when $N=1$. One can show that $\mathcal{AW}_{2}$ is a metric on $\mathcal{P}_{2}(\mathbb{R}^{N})$; its induced topology is the fundamental weak adapted topology which has several equivalent characterizations [5, 16, 26]. In this paper we focus on the case where $\mu$ and $\nu$ are Gaussian, i.e., $X$ and $Y$ are Gaussian processes.
@@ 3. The Knothe-Rosenblatt coupling
For $\mu,\nu\in\mathcal{P}(\mathbb{R}^{N})$, the Knothe-Rosenblatt coupling $\pi_{\mathrm{KR}}^{\mu,\nu}$, also called the multivariate quantile transform, defines a bicausal coupling which is known to be $\mathcal{AW}_{2}$-optimal under a stochastic co-monotonicity condition on $\mu$ and $\nu$ [9, Proposition 3.4] (also see Remark 4.7 below). See [4, 10] for recent applications of Knothe-Rosenblatt transport maps (also called monotone triangular transport maps) in statistics and machine learning.
We consider directly the Gaussian case where $\mu=\mathcal{N}(a,A)$, $\nu=\mathcal{N}(b,B)$ with $A,B\in\mathscr{S}_{++}(N)$. It is useful to parameterize $A$ and $B$ in terms of their Cholesky decompositions. Let $\mathscr{L}_{++}(N)$ (resp. $\mathscr{L}_{+}(N)$) be the space of $N$-by-$N$ lower triangular matrices with positive (resp. non-negative) diagonal entries. Then there exist unique $L,M\in\mathscr{L}_{++}(N)$ such that $A=LL^{\top}$ and $B=MM^{\top}$. It is not difficult to show (see, e.g., [21, Proposition 2]) that the Cholesky map $A\mapsto L$ is a diffeomorphism from $\mathscr{S}_{++}(N)$ onto $\mathscr{L}_{++}(N)$. When $A\in\mathscr{S}_{+}(N)$ there still exists $L\in\mathscr{L}_{+}(N)$ such that $A=LL^{\top}$, but $L$ is not necessarily unique (an example is given in (5.1)). On some probability space, let $\epsilon\sim\mathcal{N}_{N}(0,I)$ be an $N$-dimensional standard Gaussian random vector. The Knothe-Rosenblatt coupling $\pi_{\mathrm{KR}}^{\mu,\nu}$ between $\mu$ and $\nu$ is given by the law of $(X,Y)=(a+L\epsilon,b+M\epsilon)$, and is characterized by the property that each $(\pi_{\mathrm{KR}}^{\mu,\nu})_{t+1}(\cdot|x_{\mathbf{t}},y_{\mathbf{t}})$ is the comonotonic coupling between the univariate Gaussian distributions $\mu_{t+1}(\cdot|x_{\mathbf{t}})$ and $\nu_{t+1}(\cdot|y_{\mathbf{t}})$. We may also call $\pi_{\mathrm{KR}}^{\mu,\nu}$ the synchronous coupling since the same noise sequence $\epsilon$ is used to drive both $X$ and $Y$ as solutions to suitable stochastic difference equations.
For univariate processes, the Knothe-Rosenblatt coupling coincides with what we call the adapted Brenier coupling $\pi_{\mathrm{AB}}^{\mu,\nu}$ defined via (2.5) by
(3.1) $(\pi_{\mathrm{AB}}^{\mu,\nu})_{t+1}(\cdot|x_{\mathbf{t}},y_{\mathbf{t}})=\pi_{\mathrm{W}}^{\mu_{t+1}(\cdot|x_{\mathbf{t}}),\nu_{t+1}(\cdot|y_{\mathbf{t}})}.$
That is, the conditional marginals are coupled by the Brenier map which is $\mathcal{W}_{2}$-optimal. While the Knothe-Rosenblatt coupling does not have a direct analogue for multivariate processes, the property (3.1) makes sense in any dimension. In particular, if we consider (absolutely continuous) $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d\times N})$ where $d$ is the space dimension and $N$ is the time dimension, then $\pi_{\mathrm{AB}}^{\mu,\nu}=\pi_{\mathrm{KR}}^{\mu,\nu}$ when $d=1$ and $\pi_{\mathrm{AB}}^{\mu,\nu}=\pi_{\mathrm{W}}^{\mu,\nu}$ when $N=1$ (in this case $\mathcal{W}_{2}=\mathcal{AW}_{2}$). In this paper we use the terminology of Knothe-Rosenblatt coupling but note that the adapted Brenier coupling is more relevant in the study of multivariate processes. [^2] This possibility is briefly discussed in [8, Section 5].
Returning to the univariate Gaussian setting, we observe that $\pi_{\mathrm{KR}}^{\mu,\nu}$ is a deterministic coupling under which
(3.2) $Y=T_{\mathrm{KR}}^{\mu,\nu}X,\quad\text{where}\quad T_{\mathrm{KR}}^{\mu,\nu}x=b+ML^{-1}(x-a).$
Since $L,M\in\mathscr{L}_{++}(N)$, we have $ML^{-1}\in\mathscr{L}_{++}(N)$. Following [12], we define the Knothe-Rosenblatt distance $\mathcal{KR}_{2}(\mu,\nu)$ between $\mu$ and $\nu$ by
(3.3) $\mathcal{KR}_{2}^{2}(\mu,\nu)=\int_{\mathbb{R}^{N}\times\mathbb{R}^{N}}\|x-y\|^{2}\pi_{\mathrm{KR}}^{\mu,\nu}(\mathrm{d}x\mathrm{d}y).$
Since the Knothe-Rosenblatt coupling is bicausal, from (2.6) we have $\mathcal{KR}_{2}\geq\mathcal{AW}_{2}$. The following lemma is a special case of Lemma 3.3 below.
@@
For non-degenerate $\mu=\mathcal{N}(a,A=LL^{\top})$ and $\nu=\mathcal{N}(b,B=MM^{\top})$, we have $\mathcal{KR}_{2}^{2}(\mu,\nu)=\|a-b\|^{2}+d_{\mathrm{KR}}^{2}(A,B)$, where $d_{\mathrm{KR}}$ defined by
(3.4) $d_{\mathrm{KR}}^{2}(A,B)=\operatorname{tr}(A)+\operatorname{tr}(B)-2\operatorname{tr}(L^{\top}M)=\|L-M\|_{\mathrm{F}}^{2}$
is the Knothe-Rosenblatt distance and $\|L\|_{\mathrm{F}}=\sqrt{\operatorname{tr}(L^{\top}L)}$ is the Frobenius norm.
Corollary 3.2.
The Cholesky map $A\mapsto L$ is an isometry between the metric spaces $(\mathscr{S}_{++}(N),d_{\mathrm{KR}})$ and $(\mathscr{L}_{++}(N),\|\cdot\|_{\mathrm{F}})$.
The next lemma (inspired by the correlated Brownian motion in [9, 15]) provides a potential improvement of the Knothe-Rosenblatt coupling and motivates the expression of $d_{\mathrm{ABW}}$ in (1.2). We denote by $\mathscr{P}(N)$ the set of $N$-by-$N$ diagonal matrices $P=(\delta_{st}\rho_{t})$ ($\delta_{st}$ is the Kronecker delta) where $\rho_{t}\in[-1,1]$ for all $t\in[N]$.
Lemma 3.3.
Let $\mu=\mathcal{N}(a,A=LL^{\top})$ and $\nu=\mathcal{N}(b,B=MM^{\top})$ be non-degenerate. For $P\in\mathscr{P}(N)$, let $\pi^{P}$ be the law of $(X,Y)$ where
(3.5) $\begin{bmatrix}X\\ Y\end{bmatrix}\sim\mathcal{N}_{2N}\left(\begin{bmatrix}a\\ b\end{bmatrix},\begin{bmatrix}A&LPM^{\top}\\ MPL^{\top}&B\end{bmatrix}\right).$
Then $\pi^{P}\in\Pi_{bc}(\mu,\nu)$. We have $\min_{P\in\mathcal{P}}\mathbb{E}_{\pi^{P}}[\|X-Y\|^{2}]=\|a-b\|^{2}+d_{\mathrm{ABW}}^{2}(A,B)$, where $d_{\mathrm{ABW}}$ is defined by (1.2), and the minimum is attained by $P=(\delta_{st}\rho_{t})$ if and only if